the denisty of lead metal is 11.35 g/cm3 if 16.44 g of lead metal is placed in 10 ml graduated cylinder that contains 4.0 ml water what will be the final volume of water in the cylinder
1
Expert's answer
2020-11-12T07:09:44-0500

Volume of the metal = mass/density

Volume of the metal = 16.44/ 11.35 = 1.45 mL


Final volume in the cylinder = volume of the metal + volume of water

Final volume in the cylinder= 4 + 1.45 = 5.45mL